I am currently 5 times as old as my son. In 6 years time i will be three times as old as he will be then. What are our ages now
Answer by solve_for_x(190)

Let x represent your son's current age. Since you are 5 times as old as your son, your current age is 5x.

In 6 years, your son's age will be x + 6, and your age will be 5x + 6.

Since you will then be 3 times as old as your son, you can write the equation:

5x + 6 = 3(x + 6)

Solving for x gives:

5x + 6 = 3x + 18

5x - 3x + 6 = 18

2x + 6 = 18

2x + 6 - 6 = 18 - 6

2x = 12

x = 6

Solution: Your son's current age is 6, and your current age is 5(6) = 30.